Let a pair of dice be thrown and the random variable X be the sum of the numbers that appear on the two dice. Find the mean or expectation of X

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
The sample space contains 6*6=36 elements
sum of elements on dice from 2,3,4,.,12
p(sum=2) = 1/36 ;p(sum=3) = 2/36 ;p(sum=4) = 3/36
p(sum=5) = 4/36 ;p(sum=6) = 5/36 ;p(sum=7) = 6/36
p(sum=8) = 5/36 ;p(sum=9) = 4/36 ;p(sum=10) = 3/36
p(sum=11) = 2/36 ;p(sum=12) = 1/36
Hence expectation=np
=2(1/36)+3(2/36)+4(3/36)+5(4/36)+6(5/36)+7(6/36)+8(5/36)+9(4/36)+10(3/36)+11(2/36)+12(1/36)
=7
Thanks and Regards
Shaik Aasif
askIITians faculty